The magnitide of the three forces are F1 = 1.6 kN, F2 = 1.2 kN, and F3 = 1.0 kN. Compute their resultant in the form (a) R = Rxi + Ryj + Rzk; and (b) R = Rλ


1
Expert's answer
2021-06-06T16:42:52-0400

a) consider the forces being arranged as shown



In simplified form the diagram appears as shown below.


"R=2.288i+1.4056j+0.9063k"


b) "|R|=\sqrt{2.288^2+1.4056^2+0.9063^2}=2.834KN"

"R=R\lambda=2.834[\frac{2.288i+1.4056j+0.9063k]}{2.834}"

"=2.834[0.8073i+0.496j+0.32k]"